The relative luminosity of wavelength 600 nm is 0.6. Find the radiant flux of 600 nm needed to produce the same brightness sensation as produced by 120 W of radiant flux at 555 nm.

Grade:11

1 Answers

Jitender Pal
9 years ago
Sol. Relative luminousity = Lu min ous flux of a source of given wavelength/ Lu min ous flux of a source of 555 nm of same power Let the radiant flux needed be P watt. Ao, 0.6 = Lu min ous flux of source ‘P’ watt/685 P ∴ Luminous flux of the source = (685 P)× 0.6 = 120 × 685 ⇒ P = 120/0.6 = 200 W